OpenStudy (anonymous):

maclaurin series: sinh^2 (x)

OpenStudy (anonymous):

Using a power reduction formula, you have \[\sinh^2x=\frac{1}{2}\cosh2x-\frac{1}{2}\] Next, since \(\cosh x=\dfrac{e^{x}+e^{-x}}{2}\), you get \[\begin{align*}\sinh^2x&=\frac{1}{2}\left(\frac{e^{2x}+e^{-2x}}{2}\right)-\frac{1}{2}\\ &=\frac{1}{4}e^{2x}+\frac{1}{4}e^{-2x}-\frac{1}{2} \end{align*}\] Recall the power series for \(e^x\): \[e^{x}=\sum_{n=0}^\infty \frac{x^n}{n!}\] Substituting \(2x\) for \(x\), you have \[\begin{align*}\sinh^2x&=\frac{1}{4}\sum_{n=0}^\infty \frac{(2x)^n}{n!}+\frac{1}{4}\sum_{n=0}^\infty \frac{(-2x)^n}{n!}-\frac{1}{2}\\ &=\small \frac{1}{4}\left(1+2x+\frac{(2x)^2}{2!}+\frac{(2x)^3}{3!}+\cdots\right)+\frac{1}{4}\left(1-2x+\frac{(2x)^2}{2!}-\frac{(2x)^3}{3!}+\cdots\right)-\frac{1}{2}\\ &=\frac{1}{4}\left(\frac{(2x)^2}{2!}+\frac{(2x)^4}{4!}+\cdots\right)+\frac{1}{4}\left(\frac{(2x)^2}{2!}+\frac{(2x)^4}{4!}+\cdots\right)\\ &=\frac{1}{2}\left(\frac{(2x)^2}{2!}+\frac{(2x)^4}{4!}+\cdots\right)\\ &=\frac{2^1x^2}{2!}+\frac{2^3x^4}{4!}+\cdots\\ &=\large\sum_{n=1}^\infty\frac{2^{2n-1}x^{2n}}{(2n)!} \end{align*}\]